OpenStudy (anonymous):

A die is thrown 24 times. How many times you expect the number '4' to come up?

OpenStudy (anonymous):

this is based on a die with 6 faces.... So, I feel it should be (1/6) + (1/6) + (1/6) + ......... + (1/6) So, it will be 24 times (1/6) = 24*(1/6) = 4 times Pls correct me if I am wrong...

OpenStudy (anonymous):

@ganeshie8 Thanks!! ... for confirming my answer is correct....

ganeshie8 (ganeshie8):

good job !!